OpenStudy (anonymous):

The volume of a rectangular prism is equal to x 3 + 4x 2 + 3x. What is the longest edge of the prism?

OpenStudy (taylor<3srin):

x3 + 4x2 + 3x = l*w*h x*(x^2 + 4x + 3) = l*w*h x* ( x + 3) * (x + 1) = l*w*h x has to be positive --> x + 3 is the longest edge\[x3+4x2+3x=l*w*h s \]

OpenStudy (anonymous):

ok

OpenStudy (taylor<3srin):

Can I please have a medal?

OpenStudy (anonymous):

ok